OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <2, 6>

OpenStudy (jdoe0001):

take their dot product if their dot product is 0 then they're orthogonal(perpendicular)

OpenStudy (jdoe0001):

hmmm actually shoot lemme fix that

OpenStudy (jdoe0001):

parallel vectors, are the ones that have some common factor or a common scalar for example <3,4> and <9,12> are parallel because <9,12> is really just 3<3,4>

OpenStudy (anonymous):

I think I follow

OpenStudy (anonymous):

I'm still confused on how to find the dot production, mind giving me a small example?

OpenStudy (jdoe0001):

hmmm I'd assume you've covered that in the chapter by now

OpenStudy (jdoe0001):

\(\bf <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\impliedby \textit{dot product}\)

OpenStudy (anonymous):

Yea I did cover it in the lesson. So let me clarify if u is <1,5> and v was <1,3> and we simply just multiply 1*5 and 1*3. Than just add 5+3 right?

OpenStudy (jdoe0001):

hmm shoot, lemme fix that as well

OpenStudy (jdoe0001):

\(\bf <{\color{brown}{ 1,5}}>\cdot <{\color{blue}{ 1,3}}>\implies {\color{brown}{ 1}}\cdot{\color{blue}{ 1}}+{\color{brown}{ 5}}\cdot {\color{blue}{ 3}}\)

OpenStudy (anonymous):

Alright I got it. Thanks for the help!

OpenStudy (jdoe0001):

yw